# GeoVLMath: Enhancing Geometry Reasoning in Vision-Language Models via Cross-Modal Reward for Auxiliary Line Creation

### v1

@@ 2.4.2 Optimization
We adopt GRPO as the policy optimization algorithm. The overall reward signal combines the cross-modal reward introduced in Section 2.4.1 with a final-answer reward defined as a binary score, yielding 1 if the predicted final answer matches the ground truth and 0 otherwise, i.e.,
$r=\alpha r_{\text{aux}}+(1-\alpha)r_{\text{ans}}$ (3)
Given a geometric diagram $I$ and a question $q$, GRPO samples a set of response sequences $\{y_{1},y_{2},\dots,y_{G}\}$ from the old policy model $\pi_{\theta_{\text{old}}}$. The policy model $\pi_{\theta}$ is then optimized by maximizing the following objective, following the formulation introduced in (Guo et al., 2025):
$\mathcal{L}_{\text{GRPO}}=$ $\frac{1}{G}\sum_{i=1}^{G}\bigg(\min\bigg(\frac{\pi_{\theta}(y_{i}\mid I,q)}{\pi_{\theta_{\text{old}}}(y_{i}\mid I,q)}A_{i},\ \text{clip}\left(\frac{\pi_{\theta}(y_{i}\mid I,q)}{\pi_{\theta_{\text{old}}}(y_{i}\mid I,q)},\ 1-\epsilon,\ 1+\epsilon\right)A_{i}\bigg)-\beta\ \mathbb{D}_{\text{KL}}(\pi_{\theta}\parallel\pi_{\text{ref}})\bigg)$ (4)
Here, $G$ denotes the group size, while $\epsilon$ and $\beta$ are hyperparameters for clipping and KL penalty.
@@ 3 Data Creation
Building on this framework, we curate the AuxSolidMath dataset to support model training. As illustrated in Figure 3, our data creation pipeline proceeds through four progressive steps: automated problem identification, automated deduplication and diagram extraction, structured data extraction, and manual verification. The pipeline standardizes raw exam problems into semantically aligned, high-quality multimodal instances that support training vision-language models for auxiliary-line geometric reasoning (More details in Appendix D). Figure 4 shows an example from the dataset. Each instance is represented as a five-tuple consisting of the problem description, the final answer, the auxiliary-line description, the original diagram, and the auxiliary-line diagram.
Figure 3: Overview of the Proposed Data Creation Pipeline.
Automated Problem Identification. We curate AuxSolidMath from 350+ authentic high-school geometry exam sets by automatically selecting problems that explicitly require auxiliary-line constructions. A two-stage filter consists of cue-verb retrieval in reference solutions (e.g., “connect,” “construct,” “draw,” “establish”) and verification of paired diagrams (original and auxiliary-line-annotated), retaining only items that meet both criteria.
Automated Deduplication and Diagram Extraction. We automatically deduplicate the dataset and extract paired diagrams to ensure unique and high-quality instances. Problem Deduplication. Duplicates are detected by textual matching, and only the first occurrence is retained. Diagram Extraction. For each retained problem, we extract the original and auxiliary-line-annotated diagrams and apply OpenCV-based filtering to discard low-resolution or unclear images.
Structured Data Extraction. Building on the high-quality diagram pairs obtained in the previous step, we extract three textual fields for each problem: the problem description, the final answer, and the auxiliary-line description. Because the source Word files embed MathType formulas that standard parsers handle poorly, we render pages as images and parse them with a vision-language model. To ensure accuracy, we curate a 300-example seed verified by GPT-4o (Hurst et al., 2024) and use it to fine-tune Qwen2.5-VL-7B-Instruct (Bai et al., 2025), yielding a reliable, domain-adapted extractor. All outputs follow a unified JSON schema for downstream training.
Manual Verification. Each instance is independently reviewed by two AI researchers for accuracy, completeness, uniqueness, semantic consistency, as well as visual clarity and resolution. If either reviewer flags an issue, the instance is collaboratively revised, with up to three rounds per instance to systematically identify and correct even subtle or ambiguous errors. A human review resolves complex symbolic expressions and ambiguous diagrammatic elements that are often misinterpreted by automated tools. Despite its modest cost and effort, this step remains indispensable for maintaining the high data fidelity required for reliable model training and evaluation.
@@
Dataset Statistics. Our dataset AuxSolidMath comprises 3,018 solid geometry problems collected from real high school examination papers. Within this dataset, we curate a new benchmark, GeoAuxBench, designed specifically to evaluate a model’s ability to construct auxiliary lines, a skill essential to solving complex geometry problems. GeoAuxBench contains 302 examples and is divided into two difficulty levels, Easy (150) and Hard (152), using the original difficulty annotations from the source exam papers rather than post hoc labels. The tiers reflect increasing reasoning complexity along three axes: (i) the sophistication of auxiliary-line constructions, (ii) the depth of multi-step reasoning, and (iii) the degree of implicit spatial inference. This design provides a principled basis for evaluating models across different levels of geometric reasoning difficulty.
@@ 4.1 Experimental Settings
Benchmark. We evaluate on GeoAuxBench, a benchmark subset of AuxSolidMath designed to evaluate auxiliary-line constructions. Existing benchmarks largely target general geometric reasoning and seldom require the introduction of auxiliary lines, making them misaligned with our task, especially in solid geometry. GeoAuxBench spans two difficulty tiers, Easy and Hard, providing a comprehensive testbed for evaluating the geometric reasoning capabilities of LVLMs.
@@
Metrics. To evaluate model performance, we adopt the Pass@k (Chen et al., 2021b) metric, a widely used evaluation criterion originally introduced by OpenAI for assessing solution correctness under sampling. We report Pass@1 and Pass@5: Pass@1 is the accuracy of a single sample, and Pass@5 is the proportion of problems for which at least one of five samples is correct.
Models. We evaluate GeoVLMath [^3] The code will be publicly available soon at https://github.com/PersistenceForever/GeoVLMath at two model sizes: 3B and 7B, all built on the Qwen2.5-VL backbone (Qwen2.5-VL-3B/7B) (Bai et al., 2025). To comprehensively assess its performance, we benchmark it against 18 state-of-the-art LVLMs, encompassing both closed-source and open-source models. Further implementation details are shown in Appendix E.
4.2 Main Results
We evaluate GeoVLMath against 18 baseline models on the GeoAuxBench benchmark, with results presented in Table 1. We summarize three key findings: (1) GeoVLMath consistently demonstrates superior performance on GeoAuxBench. GeoVLMath-3B and GeoVLMath-7B outperform their base models, Qwen2.5-VL-3B-Instruct and Qwen2.5-VL-7B-Instruct, on pass@5. Specifically, GeoVLMath-3B achieves an absolute gain of +3.55% (from 11.61% to 15.16%), and GeoVLMath-7B improves by +10.19% (from 15.93% to 26.12%). We attribute these gains to the auxiliary-line-aware training signal. During training, a cross-modal reward model evaluates the generated auxiliary-line description against a reference auxiliary-line diagram and returns a diagram-text alignment score. This supervision guides the model to recognize when auxiliary lines are warranted, position them in appropriate locations, and exploit the resulting constraints to complete multi-step derivations. In contrast, baseline models lack this targeted alignment and typically do not proactively construct auxiliary lines, instead relying on direct theorem application and consequently overlooking latent geometric structure. As a result, GeoVLMath exhibits stronger diagram-text grounding, clearer reasoning, and more reliable solutions. (2) GeoAuxBench-Hard is a challenging benchmark that clearly distinguishes LVLM capabilities in auxiliary-line-aware geometric reasoning. On the pass@1 metric, GeoAuxBench-Hard yields a clear separation in performance: the leading models, Gemini-2.5-Flash and gpt-5-mini, attain only 63.16%, whereas the open-source Qwen2.5-VL-72B-Instruct reaches 13.16%. By design, the benchmark demands deliberate auxiliary-line construction and multi-step spatial reasoning while minimizing shortcut opportunities, and it provides reference diagrams for fine-grained error analysis. These attributes establish GeoAuxBench-Hard as a concise yet highly discriminative testbed for reliably distinguishing LVLM capabilities. (3) Model scale alone does not compensate for insufficient auxiliary-line awareness. On GeoAuxBench-Easy, GeoVLMath-7B achieves higher pass@5 than Qwen2.5-VL-32B-Instruct (35.56% vs. 23.33%). On GeoAuxBench-Hard, GeoVLMath-7B also outperforms Qwen2.5-VL-32B-Instruct (16.67% vs. 13.16%). Error analysis indicates that the Qwen2.5-VL-32B-Instruct rarely constructs auxiliary lines, thereby overlooking latent spatial constraints. Instead, GeoVLMath-7B more proactively introduces auxiliary lines in appropriate locations and exploits the resulting constraints in subsequent steps. These results indicate that supervision of the auxiliary line, rather than parameter count alone, is the decisive factor for reliable reasoning.
4.3 Cross-modal Reward Model
Leveraging AuxSolidMath triplets $\langle I,d_{\text{aux}},I^{+}\rangle$, we apply rule-based perturbations to simulate typical auxiliary-line errors and use the resulting data to train a cross-modal reward model on Qwen2.5-VL-7B. The dataset comprises 2,970 training examples and 330 test examples. We train the model for 3 epochs with a batch size of 16, using the AdamW optimizer with a learning rate of 2e-5 and a cosine learning rate scheduler with a 0.1 warm-up ratio. During training, the vision tower and projection modules are frozen, while the language model remains trainable. The model achieves a pass@1 accuracy of 98.18% on the test set, indicating reliable alignment between textual auxiliary-line descriptions and their visually annotated counterparts. This result further validates cross-modal supervision as an effective means of producing geometry-aware rewards, providing signals that assess whether the proposed constructions satisfy the underlying geometric constraints of the diagram without resorting to image manipulation or explicit coordinates.
@@ 4.3 Cross-modal Reward Model
gpt-5-mini   89.33 92.67   63.16 75.00   76.25 83.84
o4-mini-2025-04-16   84.00 93.33   60.53 74.34   72.27 83.84
GPT-4o   8.67 25.33   6.58 15.13   7.63 20.23
Gemini-2.0-Flash   37.33 62.67   25.00 39.47   31.17 51.07
Gemini-2.5-Flash   84.00 91.33   63.16 78.95   73.58 85.14
Claude 3.7 Sonnet 20250219   15.33 41.33   13.16 28.29   14.25 34.81
Claude Sonnet 4 20250514   56.00 77.33   30.92 44.74   43.46 61.04
@@
InternVL3-8B   9.33 25.33   5.92 15.79   7.63 20.56
Llama-3.2-11B-Vision-Instruct   2.00 12.00   3.29 5.92   2.65 8.96
InternVL3-14B   13.33 28.67   5.92 15.13   9.63 21.90
Qwen2.5-VL-3B-Instruct   2.00 14.89   1.97 8.33   1.99 11.61
GeoVLMath-3B (Ours)   12.89 20.44   5.70 9.87   9.30 15.16
Qwen2.5-VL-7B-Instruct   5.14 20.67   3.95 11.18   4.55 15.93
GeoVLMath-7B (Ours)   14.67 35.56   5.92 16.67   10.30 26.12
@@
Qwen2-VL-72B-Instruct   6.00 15.33   5.26 8.55   5.63 11.94
Qwen2.5-VL-32B-Instruct   20.67 23.33   11.18 13.16   15.93 18.25
Llama-4-Scout-17B-16E-Instruct   20.67 36.67   7.89 18.42   14.28 27.55
InternVL3-38B   19.33 41.33   10.53 21.71   14.93 31.52
Qwen2.5-VL-72B-Instruct   24.00 40.67   13.16 19.74   18.58 30.21
InternVL3-78B   16.67 36.67   9.21 21.05   12.94 28.86
@@
4.4 Ablation Studies
Cross-Modal Reward. We assess the role of cross-modal supervision with two variants, keeping all other settings unchanged. (a) w/o cross-modal reward. This variant removes all supervision related to auxiliary lines and trains the model solely for final-answer accuracy, with no supervision on whether auxiliary lines are introduced. This isolates the effect of answer-only supervision and approximates a setting where auxiliary lines are omitted from the training objective. (b) Textual reward. Cross-modal consistency is replaced by a text-only semantic similarity objective that measures the proximity between the generated auxiliary-line description and the ground-truth annotation. Concretely, we use EmbeddingGemma (DeepMind, 2025c) to encode sentences and compute a similarity score for training. This variant favors fluent textual descriptions but does not enforce grounding to the input diagram. Findings. As reported in Table 2, removing the cross-modal reward results in performance degradation (Average pass@1: 10.30 $\rightarrow$ 7.86, pass@5: 26.12 $\rightarrow$ 22.91), underscoring the importance of geometry-aware supervision for instances that require introducing auxiliary lines. Substituting it with a purely textual similarity objective performs even worse (Average pass@1: 10.30 $\rightarrow$ 7.53, pass@5: 26.12 $\rightarrow$ 20.47), consistent with our pilot finding in Section 1 that incorrect auxiliary lines can be worse than none. These declines suggest that lexical alignment introduces spurious signals and conflicts with precise diagram grounding, favoring surface-level paraphrases over geometry-aware reasoning. Error analysis reveals distinct failure modes: (a) often ignores auxiliary-line construction and overfits to answer-only cues; (b) produces fluent but visually inconsistent descriptions (e.g., incorrect lines) that fail to constrain diagram-based reasoning. Overall, text-only alignment is intrinsically lossy for geometric structure. Robust auxiliary-line reasoning requires visually grounded, structure-preserving diagram-text alignment that enables rigorous checks of metric accuracy and incidence relations.
Table 2: Results of ablation studies (%).
@@ 4.4 Ablation Studies
Reinforcement Learning. To quantify the contribution of reinforcement learning, we ablate it and train the SFT-only variant. As shown in Table 2, GeoVLMath-7B with SFT+RL consistently outperforms the SFT-only model, with pronounced drops once RL is removed. This improvement reflects how RL drives the policy beyond strict imitation: reward-aligned optimization not only encourages exploration of more effective strategies but also provides credit assignment for beneficial intermediate steps, rather than confining learning to surface-level matching. As a result, the model becomes less dependent on dataset-specific heuristics and exhibits more stable performance across the evaluated benchmarks. From this perspective, RL functions as a post-SFT catalyst that unlocks latent capacities of the base model and consolidates preliminary SFT competence into robust multistep reasoning, particularly in scenarios demanding deliberate auxiliary-line construction.

### v2

@@ 2.4.2 Optimization
We adopt GRPO as the policy optimization algorithm. The overall reward signal combines the cross-modal reward with a final-answer reward defined as a binary score, yielding 1 if the predicted final answer matches the ground truth and 0 otherwise, i.e., $r=\alpha r_{\text{aux}}+(1-\alpha)r_{\text{ans}}$.
Given a geometric diagram $I$ and a question $q$, GRPO samples a set of response sequences $\{y_{1},y_{2},\dots,y_{G}\}$ from the old policy $\pi_{\theta_{\text{old}}}$. The policy $\pi_{\theta}$ is then optimized by maximizing the following objective:
$\mathcal{L}_{\text{GRPO}}=$ $\frac{1}{G}\sum_{i=1}^{G}\bigg(\min\bigg(\frac{\pi_{\theta}(y_{i}\mid I,q)}{\pi_{\theta_{\text{old}}}(y_{i}\mid I,q)}A_{i},\ \text{clip}\bigg(\frac{\pi_{\theta}(y_{i}\mid I,q)}{\pi_{\theta_{\text{old}}}(y_{i}\mid I,q)},\,$ (2)
$\qquad 1-\epsilon,1+\epsilon\bigg)A_{i}\bigg)-\beta\,\mathbb{D}_{\text{KL}}(\pi_{\theta}\parallel\pi_{\text{ref}})\bigg)$
Here, $G$ denotes the group size, while $\epsilon$ and $\beta$ are hyperparameters for clipping and the KL penalty.
@@ 3 Data Creation
We curate the AuxSolidMath dataset to support training and evaluation of auxiliary-line reasoning in solid geometry. As illustrated in Figure 3, our data creation pipeline proceeds through four progressive steps: automated problem identification, automated deduplication and diagram extraction, structured data extraction, and manual verification. The pipeline standardizes raw exam problems into semantically aligned, high-quality multimodal instances that support training vision-language models for auxiliary-line geometric reasoning (More details in Appendix B). Figure 4 shows an example from the dataset. Each instance is represented as a five-tuple consisting of the problem description, the final answer, the auxiliary-line description, the original diagram, and the auxiliary-line diagram.
Automated Problem Identification. We curate AuxSolidMath from over 350 authentic high-school geometry exam sets using a two-stage filter that selects problems explicitly requiring auxiliary-line constructions, based on cue-verb retrieval from reference solutions and verification of paired original and auxiliary-line–annotated diagrams.
Automated Deduplication and Diagram Extraction. We automatically deduplicate the dataset and extract paired diagrams to ensure unique and high-quality instances. Duplicates are detected by textual matching, and only the first occurrence is retained. For each retained problem, we extract the original and auxiliary-line-annotated diagrams and apply OpenCV-based filtering to discard low-resolution or unclear images.
Structured Data Extraction. Building on the high-quality diagram pairs obtained in the previous step, we extract three textual fields for each problem: the problem description, the final answer, and the auxiliary-line description. Because the source Word files embed MathType formulas that standard parsers handle poorly, we render pages as images and parse them with a fine-tuned LVLM. All outputs are packaged in a structured JSON format.
Manual Verification. Each instance is independently reviewed by two AI researchers to ensure accuracy, completeness, uniqueness, semantic consistency, and visual quality. Instances flagged by either reviewer undergo collaborative re-examination, with up to three review rounds to identify and correct subtle or ambiguous errors. This step resolves complex symbolic expressions and diagrammatic ambiguities that automated tools often misinterpret, and remains indispensable for ensuring high data fidelity in model training and evaluation.
@@
Dataset Statistics. AuxSolidMath comprises 3,018 solid geometry problems collected from real high school examination papers. Within this dataset, we curate a new benchmark, GeoAuxBench, designed specifically to evaluate a model’s ability to construct auxiliary lines, a skill essential to solving complex geometry problems. GeoAuxBench contains 302 examples and is divided into two difficulty levels, Easy (150) and Hard (152), using the original difficulty annotations from the source exam papers rather than post hoc labels. Hard problems involve reasoning over implicit spatial relations (e.g., cross-plane or hidden projections), while Easy problems rely on explicit relations. By inheriting these exam-defined criteria, GeoAuxBench supports realistic evaluation of auxiliary-line construction in solid geometry.
@@ 4.1 Experimental Settings
Benchmark. We conduct our evaluation on GeoAuxBench, a newly introduced benchmark for auxiliary-line construction with two difficulty tiers: Easy and Hard. We further report results on widely used benchmarks, i.e., MathVision (Wang et al., 2024a), SolidGeo (Wang et al., 2025a), and the OlympiadBench subset of SolidGeo.
@@
Metrics. We evaluate model performance using Pass@k (Chen et al., 2021b), a widely adopted metric introduced by OpenAI. Specifically, we report Pass@1 and Pass@5: Pass@1 measures the accuracy of a single generated solution, while Pass@5 denotes the proportion of problems for which at least one of five generated solutions is correct.
Models. We assess  [^3] The code will be publicly available soon at https://github.com/PersistenceForever/GeoVLMath at two model scales, 3B and 7B, both built upon the Qwen2.5-VL backbone. We compare GeoVLMath against 20 strong LVLM baselines, covering both closed-source and open-source models, as well as representative methods following the code-driven visual construction paradigm, including V-Thinker Qiao et al. (2025) and CodePlot-CoT Duan et al. (2025).
@@
gpt-5-mini   89.33 92.67   63.16 75.00   76.16 83.78
o4-mini-2025-04-16   84.00 93.33   60.53 74.34   72.19 83.77
GPT-4o   8.67 25.33   6.58 15.13   7.62 20.20
Gemini-2.0-Flash   37.33 62.67   25.00 39.47   31.12 50.99
Gemini-2.5-Flash   84.00 91.33   63.16 78.95   73.51 85.10
Claude 3.7 Sonnet 20250219   15.33 41.33   13.16 28.29   14.24 34.77
Claude Sonnet 4 20250514   56.00 77.33   30.92 44.74   43.38 60.93
@@
V-Thinker (7B)   8.00 17.33   3.29 10.53   5.63 13.91
InternVL3-8B   9.33 25.33   5.92 15.79   7.61 20.53
Llama-3.2-11B-Vision-Instruct   2.00 12.00   3.29 5.92   2.65 8.94
InternVL3-14B   13.33 28.67   5.92 15.13   9.60 21.86
Qwen2.5-VL-3B-Instruct   2.00 14.89   1.97 8.33   1.98 11.59
GeoVLMath-3B (Ours)   12.89 20.44   5.70 9.87   9.27 15.12
Qwen2.5-VL-7B-Instruct   5.14 20.67   3.95 11.18   4.54 15.89
GeoVLMath-7B (Ours)   14.67 35.56   5.92 16.67   10.27 26.05
@@
CodePlot-CoT (32B)   4.00 8.67   2.63 5.26   3.31 6.95
Qwen2-VL-72B-Instruct   6.00 15.33   5.26 8.55   5.63 11.92
Qwen2.5-VL-32B-Instruct   20.67 23.33   11.18 13.16   15.89 18.21
Llama-4-Scout-17B-16E-Instruct   20.67 36.67   7.89 18.42   14.24 27.48
InternVL3-38B   19.33 41.33   10.53 21.71   14.90 31.46
Qwen2.5-VL-72B-Instruct   24.00 40.67   13.16 19.74   18.54 30.14
InternVL3-78B   16.67 36.67   9.21 21.05   12.92 28.81
@@
LVLM   MathVision   OlympiadBench   SolidGeo
  Pass@1 Pass@5   Pass@1 Pass@5   Pass@1 Pass@5
V-Thinker (7B)   16.36 31.82   4.60 13.22   10.39 29.38
CodePlot-CoT (32B)   10.00 29.09   2.87 5.17   16.62 30.86
Qwen2.5-VL-3B-Instruct   5.45 16.36   2.30 6.32   5.04 12.17
GeoVLMath-3B (Ours)   5.45 20.91   14.94 25.86   14.54 32.34
Qwen2.5-VL-7B-Instruct   11.82 20.91   2.87 13.22   5.34 14.84
GeoVLMath-7B (Ours)   16.36 36.36   25.86 50.57   17.51 42.34
Table 2: Overall evaluation on widely used benchmarks (%).
4.2 Main Results
As shown in Table 1 and Table 2, we make three key observations: (1) GeoVLMath achieves strong and robust performance among models of comparable scale across benchmarks. On GeoAuxBench, both GeoVLMath-3B and GeoVLMath-7B outperform their corresponding base models, Qwen2.5-VL-3B-Instruct and Qwen2.5-VL-7B-Instruct, on pass@5. GeoVLMath-3B improves by +3.53% (11.59% $\rightarrow$ 15.12%), while GeoVLMath-7B yields a larger gain of +10.16% (15.89% $\rightarrow$ 26.05%). Moreover, GeoVLMath consistently surpasses the representative code-driven visual construction method V-Thinker on Pass@5, improving by +12.14% (13.91% $\rightarrow$ 26.05%), despite not relying on executable drawing code or intermediate diagram rendering. Beyond this task-aligned benchmark, GeoVLMath also achieves the best performance among the compared models on widely used public benchmarks. These results demonstrate that the auxiliary-line-aware training signal generalizes beyond GeoAuxBench and improves geometric reasoning robustness. (2) GeoAuxBench-Hard is a challenging benchmark that clearly distinguishes LVLM capabilities in geometric reasoning. For example, even strong closed-source models such as Gemini-2.5-Flash and gpt-5-mini achieve Pass@1 scores of 63.16%, whereas Qwen2.5-VL-72B-Instruct attains only 13.16%, highlighting a substantial performance gap. This is primarily because the benchmark demands deliberate auxiliary-line construction and multi-step spatial reasoning while minimizing shortcut opportunities, and provides reference diagrams for fine-grained error analysis. Together, GeoAuxBench-Hard serves as a concise yet highly discriminative testbed for differentiating LVLM capabilities. (3) Model scale alone does not compensate for insufficient auxiliary-line awareness. On GeoAuxBench-Easy, GeoVLMath-7B achieves higher pass@5 than Qwen2.5-VL-32B-Instruct (35.56% vs. 23.33%). On GeoAuxBench-Hard, GeoVLMath-7B also outperforms Qwen2.5-VL-32B-Instruct (16.67% vs. 13.16%). Error analysis shows that Qwen2.5-VL-32B-Instruct rarely constructs auxiliary lines and thus overlooks latent spatial constraints, whereas GeoVLMath-7B proactively introduces appropriate auxiliary lines and exploits the induced constraints during reasoning. This comparison highlights that auxiliary-line supervision, rather than model scale alone, is critical for reliable geometric reasoning.
4.3 Cross-modal Reward Model
Leveraging AuxSolidMath triplets $\langle I,d_{\text{aux}},I^{+}\rangle$, we apply rule-based perturbations to simulate typical auxiliary-line errors and use the resulting data to train a cross-modal reward model on Qwen2.5-VL-7B. The dataset comprises 2,970 training examples and 330 test examples. We train the model for 3 epochs with a batch size of 16, using the AdamW optimizer with a learning rate of 2e-5 and a cosine learning rate scheduler with a 0.1 warm-up ratio. During training, the vision tower and projection modules are frozen, while the language model remains trainable. The model achieves a pass@1 accuracy of 98.18% on the test set, indicating reliable alignment between textual auxiliary-line descriptions and their visually annotated counterparts.
@@ 4.3 Cross-modal Reward Model
Table 3: Results of ablation studies (%).
4.4 Ablation Studies
Cross-Modal Reward. We assess the role of cross-modal supervision with two variants, keeping all other settings unchanged. (a) w/o cross-modal reward. This variant removes all supervision related to auxiliary lines and trains the model solely for final-answer accuracy, with no supervision on whether auxiliary lines are introduced. This allows us to assess the effect of training with answer-only supervision, approximating a setting where auxiliary lines are omitted from the training objective. (b) Textual reward. In this variant, cross-modal consistency is replaced with a text-only semantic similarity objective that evaluates the similarity between the generated auxiliary-line description and the ground-truth annotation. Concretely, we use EmbeddingGemma (DeepMind, 2025a) to encode sentences and compute a similarity score for training. This variant favors fluent textual descriptions but does not enforce grounding to the input diagram. Findings. As reported in Table 3, removing the cross-modal reward results in performance degradation (Average pass@1: 10.27 $\rightarrow$ 7.83, pass@5: 26.05 $\rightarrow$ 22.85), underscoring the importance of geometry-aware supervision for instances that require introducing auxiliary lines. Substituting it with a purely textual similarity objective performs even worse (Average pass@1: 10.27 $\rightarrow$ 7.51, pass@5: 26.05 $\rightarrow$ 20.42), consistent with our pilot finding in Section 1 that incorrect auxiliary lines can be worse than none. These declines suggest that lexical alignment introduces spurious signals and conflicts with precise diagram grounding, favoring surface-level paraphrases over geometry-aware reasoning. Error analysis reveals distinct failure modes: (a) often ignores auxiliary-line construction and overfits to answer-only cues; (b) produces fluent but visually inconsistent descriptions (e.g., incorrect lines) that fail to constrain diagram-based reasoning. Overall, text-only alignment fails to faithfully capture geometric structure, motivating the need for visually grounded, structure-preserving diagram-text alignment to support rigorous verification of geometric measurements and structural relations.
Reinforcement Learning. To quantify the contribution of reinforcement learning, we remove the RL stage and train an SFT-only variant. As shown in Table 3, GeoVLMath-7B trained with SFT+RL consistently outperforms its SFT-only counterpart, with clear performance degradation when the RL stage is removed. This improvement reflects the role of RL in moving beyond strict imitation: reward-aligned optimization encourages exploration of more effective strategies and enables credit assignment for beneficial intermediate steps, rather than relying on surface-level matching alone. As a result, RL acts as a post-SFT catalyst that consolidates preliminary SFT competence into robust multi-step reasoning, particularly in scenarios requiring auxiliary-line construction.